# Towards Tracing Trustworthiness Dynamics: Revisiting Pre-training Period of Large Language Models

### v1

@@ 4 Probing LLMs using Mutual Information
Recently, Choi et al. (2023) shows that mutual information estimation is bounded by linear probing accuracy. Also, the mutual information can be used to investigate the dynamics of neural networks during training (Shwartz-Ziv and Tishby, 2017; Saxe et al., 2019; Goldfeld and Polyanskiy, 2020; Pimentel et al., 2020; Geiger, 2021; Lorenzen et al., 2021; Zhou et al., 2023). Therefore, motivated by the above, we adopt a different perspective by probing LLM checkpoints through the lens of mutual information, particularly focusing on the aforementioned trustworthiness dimensions.
We explain our probing strategy and experimental setup in Section 4.1 and Section 4.2, respectively. The empirical observations are shown and analyzed in Section 4.3. In particular, we find that there is a phase transition from “fitting” to “compression” during the pre-training period of LLMs, which is consistent with previous study on traditional DNNs (Shwartz-Ziv and Tishby, 2017; Noshad et al., 2019).
@@ 4.1 Probing Strategy
Figure 6: The dynamics of $I(T,X)$ and $I(T,Y)$ for TruthfulQA across various layers during pre-training. The similar trend in other datasets is in Appendix E.2.
@@ 4.2 Experimental Setup
Following the official code and reported hyperparameters from Liu et al. (2023e), we initiate pre-training from a randomly initialized model using the corpus for the first checkpoint, and save more granular checkpoints to observe finer experimental phenomena. More discussions are available in Appendix C.
@@ 4.3 The Dynamics of Pre-training
Figure 6 shows that $I(T,X)$ generally exhibits an initial increase followed by a decrease across all the considered layers during pre-training. And $I(T,Y)$ continues to show a consistent upward trend. Note that middle layer representations exhibit a larger $I(T,Y)$ comparing to that from other layers. It suggests that middle layer representations encode more information about the opposing concepts of trustworthiness.
@@
Overall, considering $I(T,X)$ and $I(T,Y)$ collectively, it becomes evident that there are two phases during pre-training. In the first and shorter phase, both $I(T,X)$ and $I(T,Y)$ increase. While in the second and much longer phase, $I(T,X)$ decreases and $I(T,Y)$ continues to increase. Although our strategy is completely different from Shwartz-Ziv and Tishby (2017), the two-phase phenomenon exhibits similarities. At the beginning of pre-training, the randomly-initialized LLM fails to preserve the relevant information, so $I(T,X)\approx 0$ and $I(T,Y)\approx 0$. Next, as LLM gradually fits the pre-training dataset, its abilities in language understanding and concept modeling enhance, contributing to increases in both $I(T,X)$ and $I(T,Y)$. As pre-training progresses, LLM learns to better compress the irrelevant information in the dataset and preserve more label-related information (i.e., trustworthiness), leading to a reduction in $I(T,X)$ and an improvement in $I(T,Y)$. Overall, we are at the forefront of investigating the phase transition from “fitting” to “compression” in the context of trustworthiness during pre-training. It is our hope that our insights will motivate further exploration into the pre-training dynamics of LLMs.
@@ 5 Related Work
Probing classifiers (Alain and Bengio, 2016; Tenney et al., 2019; Pimentel et al., 2020; Li et al., 2021; Belinkov, 2022; Räuker et al., 2023) is one of the prominent methods for identify certain properties acquired by the language model (Zhao et al., 2023). Researchers probe LLMs and discover linear separable patterns within LLMs, including space and time (Gurnee and Tegmark, 2023), game states (Nanda et al., 2023), answerability (Slobodkin et al., 2023), and some counterfactual pairs of concepts (Park et al., 2023). It is also observed that LLM representations contain linearly separable patterns about trustworthiness, such as truthfulness (Li et al., 2023; Marks and Tegmark, 2023; Zou et al., 2023). However, they do not probe LLM representations during pre-training. In this work, we consider the whole pre-training period of LLMs and probe their presentations dynamically.
@@
To ensure the safety and trustworthiness of LLMs, some promising approaches explore the latent space, utilizing representations to improve model performance Liu et al. (2023c); Jorgensen et al. (2023). Various studies investigate activation engineering within LLMs from both theoretical and practical perspectives, affecting model performance by manipulating the model’s representational space Park et al. (2023); Turner et al. (2023); Zou et al. (2023). Furthermore, Wang and Shu (2023), Rimsky et al. (2023) and Wang et al. (2024) construct directional vectors to explore the model’s safety and alignment, with the goal of making models helpful, honest, and harmless. However, there has no investigation into how representations change during the pre-training phase of LLMs. In this paper, we explore and leverage representations during this phase, paving the way for new research avenues in activation engineering.
@@
Many empirical studies observe that DNNs tend to learn simple concepts during the learning process Arpit et al. (2017); Liu et al. (2021); Mangalam and Prabhu (2019). Furthermore, Xu et al. (2019), Liu et al. (2023a), and Tian et al. (2023) theoretically explain the learning preference of DNNs. Meanwhile, many researchers focus on analyzing the utility of fine-tuning for language models Merchant et al. (2020); Hao et al. (2020); Aghajanyan et al. (2021); Zhou and Srikumar (2022); Mosbach et al. (2020) and attempt to understand the in-context learning Ren et al. (2024). However, few previous studies investigate how trustworthiness is learned by LLMs during pre-training. In this paper, we take a closer look at the learning dynamic of trustworthiness within LLMs’ representations.
@@ 6 Conclusion
In this work, we take an initial and illuminating step towards elucidating the conceptual understanding of trustworthiness during pre-training. Firstly, by linear probing LLMs across reliability, privacy, toxicity, fairness, and robustness, we investigate the ability of LLMs representations to discern opposing concepts within each trustworthiness dimension during the whole pre-training period. Furthermore, motivated by the probing results, we conduct extensive experiments to reveal the potential of utilizing representations from LLMs during its previous pre-training period to enhance LLMs’ own trustworthiness. Finally, we use mutual information to probe LLMs during pre-training and reveal some similarity of the learning mechanism between LLMs and traditional DNNs. Taken collectively, the empirical study presented in this work can not only justify the potential to improve trustworthiness of LLMs using their own pre-training checkpoints, but may also lead to a better understanding of the dynamics of LLM representations, especially the trustworthiness-related concepts.
7 Limitations
There are several limitations of this work. Firstly, we only focus on five essential trustworthiness dimensions and do not encompass all the dimensions, such as those appeared in Commission et al. (2019); Liu et al. (2023b). And a wide variety of definitions for each trustworthiness dimension, as discussed by Wang et al. (2023a); Sun et al. (2024), are not completely covered in our analysis. Secondly, due to limitations in computational resources as well as the lack of open-source pre-training LLM checkpoints, we only conduct experiments on LLM360 (Liu et al., 2023e). Finally, for evaluation of TruthfulQA, the precision of evaluation results depends on the performance of the “GPT-judge” evaluator. And for multiple-choice evaluation, the evaluation results may rely on the instruction following ability of LLMs.
8 Broader Impact and Ethics Statement
This study concentrates on better understanding the learning dynamics of LLM trustworthiness during pre-training. The motivation of our steering vector experiments is centered on improving the trustworthiness of LLMs. We recognize the sensitive nature of our research and assure that it strictly complies with legal and ethical guidelines.
@@ Appendix A Guidelines for Trustworthy LLMs
The surge of LLMs brings significant concerns regarding their trustworthiness, which pertains to the aspects and extent to which humans can trust AI. Existing research in AI governance and trustworthy LLMs provides a guidance for establishing a comprehensive and reliable dimensions of trustworthy LLMs in this study.
@@ Appendix B Datasets of Truthworthy LLMs
Considering five aspects of trustworthiness: reliability, toxicity, privacy, fairness and robustness, we carefully design five binary NLP datasets. These datasets are tailored from independent lines of trustworthy AI research, with labels indicating whether a sentence satisfies each aforementioned aspect of trustworthiness. In other words, the label indicates whether the corresponding sentence contains untrue (or unfair, toxic, privacy-leakaging and perturbed) information.
The datasets considered below are balanced, i.e., the number of positive and negative numbers are almost the same. In other words, some special case, for example, the random classifier on these datasets, will achieve an accuracy around 50%.
@@
We use StereoSet (Nadeem et al., 2021) to measure the stereotype modeling ability of LLMs, i.e., whether LLMs capture stereotypical biases about race, religion, profession, and gender. Taking inter-sentence tests as the original dataset, we concatenate the context and the candidate sentence into one sentence, and the corresponding class label follows the candidate sentences capturing stereotypical, anti-stereotypical, and unrelated associations. We assign a binary label to every sentence to indicate whether it contains stereotypical bias.
@@
Following the construction of AdvGLUE benchmark (Wang et al., 2021), we perturb GLUE benchmark (Wang et al., 2018) in a human-imperceptible way. Specifically, we introduce typos by randomly change the case of 20% letters in each sentence from SST-2 (Socher et al., 2013) validation set. We assign a binary label to every sentence to indicate whether it has been attacked.
@@ Appendix C More Detailed Experimental Settings
Dimention
Reliability
Toxicity
Privacy
Fairness
Robustness
Benchmark
TruthfulQA
ToxiGen
ConfAIde
StereoSet
SST-2
Evaluation Metrics
Truthand Info%
Toxic Ratio
Accuracy
Accuracy
Accuracy
The meaning of labels in activation datasets
$y=0$: statements with false answer
@@
Within each dataset, following Li et al. (2023), we first split the original dataset into a development set and a test set at a 1:1 ratio. We further divide the development set into a training/validation set at a 4:1 ratio for the training and evaluation of the linear probe, with the steering vector also being constructed based on the development set. The test set is used to assess model performance, ensuring no data leakage occurs during the experiment.
@@
For TruthfulQA, we adopt the QA prompts following InstructGPT  Ouyang et al. (2022). Additionally, two fine-tuned GPT-3 models, i.e. a “GPT-judge”ft:davinci-002:zy-pj-035:truthfulqa-truth:8nKPYSTt and a “GPT-info,”ft:davinci-002:zy-pj-035:truthfulqa-info:8nJbtN57 are used to predict the truthfulness and informativeness of the generated outputs from LLMs, respectively. For ToxiGen, we follow Touvron et al. (2023b), employing the default ToxiGen classifier Hartvigsen et al. (2022) fine-tuned on RoBERTa Liu et al. (2020) to evaluate the toxicity of contents generated by LLMs, and finally reporting the proportion of generated text classified as toxic. For ConfAIde, we use the the tier 2 task to assess the agreement on privacy information usage. We employ the same evaluation prompt as ConfAIde Mireshghallah et al. (2023), with the adaptation of converting multiple-choice questions into binary classification tasks to evaluate the accuracy. For StereoSet, following TrustLLM (Sun et al., 2024), we provide prompts using the same template for stereotype recognition task as theirs. The generated choices are then compared with the ground-truth labels to obtain the accuracy. For perturbed SST-2, we follow Wang et al. (2023b) and use the same prompt as theirs. TruthfulQA is evaluated in a 6-shot setting, whereas other benchmarks are conducted with 0-shot settings.
@@
Regarding perplexity, we follow Radford et al. (2019) to calculate the perplexity on LAMBADA Paperno et al. (2016). The perplexity value reported for GPT-2 in Radford et al. (2019) is 8.6, and the perplexity we tested for AmberChat is 4.5. Based on our observations, we consider a perplexity value of less than 6 to be a reasonable threshold, please refer to Appendix G for examples.
@@
In our initial experimental observations using the pre-training checkpoints released in Liu et al. (2023e), we noticed that the mutual information $I(T,X)$ appeared to be consistently decreasing, which contradicts the existing two-phase phenomenon Shwartz-Ziv and Tishby (2017). This led us to speculate the possibility of overlooked experimental insights between the initial model state and the first checkpoint. Therefore, to observe more finer-grained dynamics during the pre-training phase, we utilized the official code released by Liu et al. (2023e)https://github.com/LLM360/amber-train, ensuring the hyperparameters are consistent with those reported in the original paper. We initiate pre-training from a randomly initialized model using the corpus for the first checkpoint, and saved more finely-grained checkpoints to observe finer experimental phenomena.

### v2

@@ 4 Probing LLMs using Mutual Information
Recently, Choi et al. (2023) shows that mutual information estimation is bounded by linear probing accuracy. Also, the mutual information can be used to investigate the dynamics of neural networks during training (Shwartz-Ziv and Tishby, 2017; Saxe et al., 2019; Goldfeld and Polyanskiy, 2020; Pimentel et al., 2020; Geiger, 2021; Lorenzen et al., 2021; Zhou et al., 2023b). Therefore, motivated by the above, we adopt a different perspective by probing LLM checkpoints through the lens of mutual information, particularly focusing on the aforementioned trustworthiness dimensions.
We explain our probing strategy and experimental setup in Section 4.1 and Section 4.2, respectively. The empirical observations are shown and analyzed in Section 4.3. In particular, we find that there is a phase transition from “fitting” to “compression” during the pre-training period of LLMs, which is consistent with previous studies on traditional DNNs (Shwartz-Ziv and Tishby, 2017; Noshad et al., 2019).
@@ 4.1 Probing Strategy
Figure 6: The dynamics of $I(T,X)$ and $I(T,Y)$ for TruthfulQA across various layers during pre-training. A similar trend in other datasets is in Appendix E.2.
@@ 4.2 Experimental Setup
Following the official code and reported hyperparameters from Liu et al. (2023e), we initiate pre-training from a randomly initialized model using the corpus for the first checkpoint, and save more granular checkpoints to observe finer phenomena. More discussions are available in Appendix C.
@@ 4.3 The Dynamics of Pre-training
Figure 6 shows that $I(T,X)$ generally exhibits an initial increase followed by a decrease across all the considered layers during pre-training. And $I(T,Y)$ continues to show a consistent upward trend. Note that middle layer representations exhibit a larger $I(T,Y)$ compared to that from other layers. It suggests that middle-layer representations encode more information about the opposing concepts of trustworthiness.
@@
Overall, considering $I(T,X)$ and $I(T,Y)$ collectively, it becomes evident that there are two phases during pre-training. In the first and shorter phase, both $I(T,X)$ and $I(T,Y)$ increase. While in the second and much longer phase, $I(T,X)$ decreases and $I(T,Y)$ continues to increase. Although our strategy is completely different from Shwartz-Ziv and Tishby (2017), the two-phase phenomenon exhibits similarities. At the beginning of pre-training, the randomly initialized LLM fails to preserve the relevant information, so $I(T,X)\approx 0$ and $I(T,Y)\approx 0$. Next, as LLM gradually fits the pre-training dataset, its abilities in language understanding and concept modeling enhance, contributing to increases in both $I(T,X)$ and $I(T,Y)$. As pre-training progresses, LLM learns to better compress the irrelevant information in the dataset and preserve more label-related information (i.e., trustworthiness), leading to a reduction in $I(T,X)$ and an improvement in $I(T,Y)$. Overall, we are at the forefront of investigating the phase transition from “fitting” to “compression” in the context of trustworthiness during pre-training. We hope that our insights will motivate further exploration of LLMs’ pre-training dynamics.
@@ 5 Related Work
Probing classifiers (Alain and Bengio, 2016; Tenney et al., 2019; Pimentel et al., 2020; Li et al., 2021; Belinkov, 2022; Räuker et al., 2023) is one of the prominent methods for identifying certain properties acquired by the language model (Zhao et al., 2023). Researchers probe LLMs and discover linear separable patterns within LLMs, including space and time (Gurnee and Tegmark, 2023), game states (Nanda et al., 2023), answerability (Slobodkin et al., 2023), and some counterfactual pairs of concepts (Park et al., 2023). It is also observed that LLM representations contain linearly separable patterns about trustworthiness, such as truthfulness (Li et al., 2023a; Marks and Tegmark, 2023; Zou et al., 2023). However, they do not probe LLM representations during pre-training. In this work, we consider the whole pre-training period of LLMs and probe their presentations dynamically.
@@
Numerous intriguing approaches have been proposed to investigate the trustworthiness of LLMs Ouyang et al. (2022); Rafailov et al. (2024); Zhang et al. (2024a); Li et al. (2024); Ren et al. (2024b). Specifically, some promising approaches explore the latent space, utilizing representations to improve model performance Liu et al. (2023c); Jorgensen et al. (2023). Various studies investigate activation engineering within LLMs from both theoretical and practical perspectives, affecting model performance by manipulating the model’s representational space Park et al. (2023); Turner et al. (2023); Zou et al. (2023). Furthermore, Wang and Shu (2023), Rimsky et al. (2023) and Wang et al. (2024) construct directional vectors to explore the model’s safety and alignment, with the goal of making models helpful, honest, and harmless. However, there has been no investigation into how representations change during the pre-training phase of LLMs. In this paper, we explore and leverage representations during this phase, paving the way for new research avenues in activation engineering.
@@
Many empirical studies observe that DNNs tend to learn simple concepts during the learning process Arpit et al. (2017); Liu et al. (2021); Mangalam and Prabhu (2019). Furthermore, Xu et al. (2019), Liu et al. (2023a), Zhou et al. (2024), and Tian et al. (2023) theoretically explain the learning preference of DNNs. Meanwhile, many researchers focus on analyzing the utility of fine-tuning for language models Merchant et al. (2020); Hao et al. (2020); Aghajanyan et al. (2021); Zhou and Srikumar (2022); Mosbach et al. (2020) and attempt to understand the in-context learning Ren et al. (2024a). However, few previous studies investigate how trustworthiness is learned by LLMs during pre-training. In this paper, we take a closer look at the learning dynamic of trustworthiness within LLMs’ representations.
6 Discussion
As the capabilities of LLMs have increased, conventional alignment techniques that rely on “human feedback” (like RLHF) may no longer work when trying to align models that are more powerful than humans (Burns et al., 2023; Yuan et al., 2024). To address this challenge, research institutions are actively exploring new solutions. For example, OpenAI introduces “superalignment” https://openai.com/index/introducing-superalignment/ and proposes a “weak-to-strong supervision” approach (Burns et al., 2023). Also, Meta proposes a “self-reward” mechanism (Yuan et al., 2024). At the same time, more and more research focuses on the emerging field of “self-alignment” (Sun et al., 2023; Li et al., 2023b). In this paper, we provide a deeper understanding of the pre-training dynamics and successfully align the SFT model using its own pre-training checkpoints. We believe that the pre-training period is worth being explored and it may be a promising source for self-alignment.
On the other hand, to make LLMs trustworthy, recent conventional alignment methods, such as SFT and RLHF, incur high costs due to exhaustive human annotations (Wang et al., 2022; Honovich et al., 2022; Sun et al., 2023) and time-consuming instruction tuning (Zhou et al., 2023a; Chen et al., 2023a, b). In this paper, we delve into the pre-training period to enhance trustworthiness without collecting data or tuning the model. We expect more alignment approaches inspired by the pre-training phase (like Korbak et al. (2023)) and to circumvent potential alignment costs in the future.
@@ 7 Conclusion
In this work, we take an initial and illuminating step towards elucidating the conceptual understanding of trustworthiness during pre-training. Firstly, by linear probing LLMs across reliability, privacy, toxicity, fairness, and robustness, we investigate the ability of LLMs representations to discern opposing concepts within each trustworthiness dimension during the whole pre-training period. Furthermore, motivated by the probing results, we conduct extensive experiments to reveal the potential of utilizing representations from LLMs during its previous pre-training period to enhance LLMs’ own trustworthiness. Finally, we use mutual information to probe LLMs during pre-training and reveal some similarities in the learning mechanism between LLMs and traditional DNNs. Taken collectively, the empirical study presented in this work can not only justify the potential to improve the trustworthiness of LLMs using their own pre-training checkpoints but may also lead to a better understanding of the dynamics of LLM representations, especially the trustworthiness-related concepts.
Acknowledgements
We thank the anonymous reviewers for their constructive suggestions to improve the quality of this paper. This work is supported by the Beijing Natural Science Foundation (No.4222029); the National Natural Science Foundation of China (NO.62076234); the National Key Research and Development Project (No.2022YFB2703102); the “Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China”; the Beijing Outstanding Young Scientist Program (NO.BJJWZYJH012019100020098); the Public Computing Cloud, Renmin University of China; the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China (NO.2021030199); the Huawei-Renmin University joint program on Information Retrieval; and the Unicom Innovation Ecological Cooperation Plan.
Limitations
There are several limitations of this work. Firstly, we only focus on five essential trustworthiness dimensions and do not encompass all the dimensions, such as those that appeared in Commission et al. (2019); Liu et al. (2023b). A wide variety of definitions for each trustworthiness dimension, as discussed by Wang et al. (2023a); Sun et al. (2024), are not completely covered in our analysis. Secondly, due to the absence of publicly available larger and more complex LLMs (such as 13B or others) that release pre-training period checkpoints, we are limited to conducting experiments on 7B series models (we also provide some experimental results on another 7B-size model named OLMo (Groeneveld et al., 2024) in Appendix.F). Finally, for evaluation of TruthfulQA, the precision of evaluation results depends on the performance of the “GPT-judge” evaluator. And for multiple-choice evaluation, the evaluation results may rely on the instruction following ability of LLMs.
Broader Impact and Ethics Statement
This study concentrates on better understanding the learning dynamics of LLM trustworthiness during pre-training. The motivation of our steering vector experiments is centered on improving the trustworthiness of LLMs. We recognize the sensitive nature of our research and ensure that it strictly complies with legal and ethical guidelines.
@@ Appendix A Guidelines for Trustworthy LLMs
The surge of LLMs brings significant concerns regarding their trustworthiness, especially considering the security risks inherent in the models themselves and the agents based on these models Wang and Shu (2023); Ji et al. (2023); Newman (2023); Tabassi (2023); Li et al. (2024); Zhang et al. (2024a), which pertains to the aspects and extent to which humans can trust AI. Existing research in AI governance and trustworthy LLMs provides guidance for establishing comprehensive and reliable dimensions of trustworthy LLMs in this study.
@@ Appendix B Datasets of Truthworthy LLMs
Considering five aspects of trustworthiness: reliability, toxicity, privacy, fairness, and robustness, we carefully design five binary NLP datasets. These datasets are tailored from independent lines of trustworthy AI research, with labels indicating whether a sentence satisfies each aforementioned aspect of trustworthiness. In other words, the label indicates whether the corresponding sentence contains untrue (or unfair, toxic, privacy-leakage, and perturbed) information.
The datasets considered below are balanced, i.e., the number of positive and negative numbers are almost the same. In other words, some special cases, for example, the random classifier on these datasets, will achieve an accuracy of around 50%.
@@
We use StereoSet (Nadeem et al., 2021) to measure the stereotype modeling ability of LLMs, i.e., whether LLMs capture stereotypical biases about race, religion, profession, and gender. Taking inter-sentence tests as the original dataset, we concatenate the context and the candidate sentence into one sentence, and the corresponding class label follows the candidate sentences, capturing stereotypical, anti-stereotypical, and unrelated associations. We assign a binary label to every sentence to indicate whether it contains stereotypical bias.
@@
Following the construction of AdvGLUE benchmark (Wang et al., 2021), we perturb GLUE benchmark (Wang et al., 2018) in a human-imperceptible way. Specifically, we select SST-2 (Socher et al., 2013) from GLUE. It is a popular dataset in robustness literature (Zhu et al., 2023; Zhang et al., 2022, 2024b; Chen et al., 2024). We introduce typos by randomly changing the case of 20% letters in each sentence from the SST-2 (Socher et al., 2013) validation set. We assign a binary label to every sentence to indicate whether it has been attacked.
@@ Appendix C More Detailed Experimental Settings
Dimension Reliability Toxicity Privacy Fairness Robustness
Benchmark TruthfulQA ToxiGen ConfAIde StereoSet SST-2
Evaluation Metrics Truthand InfoToxic Ratio Accuracy Accuracy Accuracy
The meaning of labels in activation datasets $y=0$: statements with false answer
@@
Within each dataset, following Li et al. (2023a), we first split the original dataset into a development set and a test set at a 1:1 ratio. We further divide the development set into a training/validation set at a 4:1 ratio for the training and evaluation of the linear probe, with the steering vector also being constructed based on the development set. The test set is used to assess model performance, ensuring no data leakage occurs during the experiment.
@@
For TruthfulQA, we adopt the QA prompts following InstructGPT  Ouyang et al. (2022). Additionally, two fine-tuned GPT-3 models, i.e. a “GPT-judge”ft:davinci-002:zy-pj-035:truthfulqa-truth:8nKPYSTt and a “GPT-info,”ft:davinci-002:zy-pj-035:truthfulqa-info:8nJbtN57 are used to predict the truthfulness and informativeness of the generated outputs from LLMs, respectively. For ToxiGen, we follow Touvron et al. (2023b), employing the default ToxiGen classifier Hartvigsen et al. (2022) fine-tuned on RoBERTa Liu et al. (2020) to evaluate the toxicity of contents generated by LLMs, and finally reporting the proportion of generated text classified as toxic. For ConfAIde, we use the tier 2 task to assess the agreement on privacy information usage. We employ the same evaluation prompt as ConfAIde Mireshghallah et al. (2023), with the adaptation of converting multiple-choice questions into binary classification tasks to evaluate the accuracy. For StereoSet, following TrustLLM (Sun et al., 2024), we provide prompts using the same template for the stereotype recognition task as theirs. The generated choices are then compared with the ground-truth labels to obtain accuracy. For perturbed SST-2, we follow Wang et al. (2023b) and use the same prompt as theirs. TruthfulQA is evaluated in a 6-shot setting, whereas other benchmarks are conducted in 0-shot settings.
@@
Regarding perplexity, we follow Radford et al. (2019) to calculate the perplexity on LAMBADA Paperno et al. (2016). The perplexity value reported for GPT-2 in Radford et al. (2019) is 8.6, and the perplexity we tested for AmberChat is 4.5. Based on our observations, we consider a perplexity value of less than 6 to be a reasonable threshold, please refer to Appendix H for examples.
@@
In our initial experimental observations using the pre-training checkpoints released in Liu et al. (2023e), we noticed that the mutual information $I(T,X)$ appeared to be consistently decreasing, which contradicts the existing two-phase phenomenon Shwartz-Ziv and Tishby (2017). This led us to speculate the possibility of overlooked experimental insights between the initial model state and the first checkpoint. Therefore, to observe more finer-grained dynamics during the pre-training phase, we utilized the official code released by Liu et al. (2023e)https://github.com/LLM360/amber-train, ensuring the hyperparameters are consistent with those reported in the original paper. We initiated pre-training from a randomly initialized model using the corpus for the first checkpoint and saved more finely-grained checkpoints to observe finer experimental phenomena.